# Testing the Shock-cooling Emission Model from Star-disk Collisions for Quasi-periodic Eruptions

### v3

@@ VI Conclusion
1) Within our sample, the eruption luminosity and duration constraints require that $r_{\star}\sim 3-10$ for GSN 069, $r_{\star}\sim 0.3-1$ for eRO-QPE2, $r_{\star}\sim 0.5-4$ for eRO-QPE3 and $r_{\star}\sim 2-5$ for eRO-QPE4, respectively. There is no allowed parameter space for the rest (RX J1301, eRO-QPE1, ASASSN-14ko and J0230) of the sample.
2) Introducing the full or partial tidal disruption limit further squeezes the allowed parameter space of eRO-QPE4 to $r_{\star}\sim 2$. The best candidate source eRO-QPE3 satisfies this limit well. For GSN 069 and eRO-QPE2, however, they would have suffer significant partial disruption and the accretion of the stripped mass might be the dominant energy source powering the eruptions, thus undermining the foundation of the shock-cooling emission model from star-disk collisions.
3) The influence of the orbital inclination $i$ is the most notable when the stellar orbit is extremely retrograde. In that optimistic case, a star with a small $r_{\star}$ of 0.010.10.01-0.1 would be able to reproduce the observed $t_{\rm{e}}$ and $L_{\rm{p}}$ for four sources (i.e., GSN 069, eRO-QPE1, eRO-QPE2 and eRO-QPE3). However, it is rare for the syste to have $i\sim\pi$.
@@
Our findings pose serious challenges to the shock-cooling emission model from star-disk collisions. Specifically, we find that for most of the sample, the requirements of $r_{\star}$ to reproduce both $L_{p}$ and $t_{e}$, and the requirement that the star should not be partially disrupted unless including the inclination effect, are in tension with each other. This issue is largely alleviated when the star is accompanied by a debris stream shed from the star and the stream-disk collisions produces the eruptions.
We thank an anonymous reviewer for useful comments and suggestions that helped improve the manuscript. This work is supported by the National Natural Science Foundation of China grant 12393814.

### v4

@@ VI Conclusion
1) Within our sample, the eruption luminosity and duration constraints require that $r_{\star}\sim 3-10$ for GSN 069, $r_{\star}\sim 0.3-1$ for eRO-QPE2, $r_{\star}\sim 0.5-4$ for eRO-QPE3, and $r_{\star}\sim 2-5$ for eRO-QPE4. There is no allowed parameter space for the rest (RX J1301, eRO-QPE1, ASASSN-14ko, and J0230) of the sample.
2) Introducing the full or partial tidal disruption limit further squeezes the allowed parameter space of eRO-QPE4 to $r_{\star}\sim 2$. The best candidate source eRO-QPE3 satisfies this limit well. For GSN 069 and eRO-QPE2, however, they would have suffered significant partial disruption, and the accretion of the stripped mass might be the dominant energy source powering the eruptions, thus undermining the foundation of the shock-cooling emission model from star-disk collisions.
3) The influence of the orbital inclination $i$ is the most notable when the stellar orbit is extremely retrograde. In that optimistic case, a star with a small $r_{\star}$ of 0.010.10.01-0.1 would be able to reproduce the observed $t_{\rm{e}}$ and $L_{\rm{p}}$ for four sources (i.e., GSN 069, eRO-QPE1, eRO-QPE2 and eRO-QPE3). However, it is rare for the system to have $i\sim\pi$.
@@
Our findings pose serious challenges to the shock-cooling emission model from star-disk collisions. Specifically, we find that for most of the sample, the requirements of $r_{\star}$ to reproduce both $L_{p}$ and $t_{e}$, and the requirement that the star should not be partially disrupted unless including the inclination effect, are in tension with each other. This issue is largely alleviated when the star is accompanied by a debris stream shed from the star, and the stream-disk collisions produce the eruptions.
We thank an anonymous reviewer for the useful comments and suggestions that helped improve the manuscript, and are grateful to the organizers and speakers of the TDE Full-process Orbital to Radiative Unified Modeling for the insightful discussions. This work is supported by the National Natural Science Foundation of China grant 12393814.